# On anomalous dissipation induced by transport noise

### v2

@@ A.2. Stochastic Meyers’ estimates for the turbulent Stokes system
In this subsection, prove the Meyers’ estimates for the turbulent Stokes system on $\mathbb{T}^{d}$, i.e.,
(A.12) $\partial_{t}v&=\kappa\Delta v+Q_{\xi}v+f+\sum_{n\geq 1}\big{(}\mathbb{P}\big{[}(\xi_{n}\cdot\nabla)v\big{]}+g_{n}\big{)}\,\dot{W}_{t}^{n}&\text{ on }&\mathbb{T}^{d},\\ v(0,\cdot)&=0&\text{ on }&\mathbb{T}^{d},\end{aligned}\right.$
@@
Here and below, we use the notation introduced at the beginning of Section 5. Although, it holds that $(\mathbb{H}^{1})^{*}=\mathbb{H}^{-1}$ here we will not employ such identification as it is important to keep track of constants. Note that $Q_{\xi}$ corresponds to the Itô-correction for the Stratonovich formulation of the transport noise for the turbulent Stokes system, see Section 5 and [9, Section 1].
Strong solutions to (A.12) can be defined similarly to the one of (A.1). The following complements [9, Theorem 3.2] in case of $L^{\infty}$-transport noise.
@@
Let $\kappa>0$. Assume that $\xi_{n}=(\xi_{j}^{n})_{j=1}^{d}:{\mathbb{R}}_{+}\times\Omega\times\mathbb{T}^{d}\to{\mathbb{R}}^{d}$ is $\mathscr{P}\otimes\mathscr{B}(\mathbb{T}^{d})$-measurable for all $n\geq 1$ for which there exist $M_{0}>0$ and $\kappa_{0}\in[0,\kappa)$ such that, a.e. on +×${\mathbb{R}}_{+}\times\Omega\times\mathbb{T}^{d}$,
[…]
Then there exists $p_{0}(d,\kappa,\kappa_{0},M_{0})>2$ such that, for all $T\in(0,\infty)$, $p\in[2,p_{0}]$, $q\in[2,p]$ and progressively measurable processes $f$ and $g$ satisfying
$f\in L^{p}((0,T)\times\Omega;\mathbb{H}^{-1,q}(\mathbb{T}^{d}))\quad\text{ and}\quad g=(g_{n})_{n\geq 1}\in L^{p}((0,T)\times\Omega;\mathbb{L}^{q}(\mathbb{T}^{d};\ell^{2})),$
there exists a unique strong solution $v\in L^{p}((0,T)\times\Omega;\mathbb{H}^{1,q}(\mathbb{T}^{d}))$ to (A.12) and
(A.15) $\|v\|_{L^{p}((0,T)\times\Omega;H^{1,q}(\mathbb{T}^{d};{\mathbb{R}}^{d}))}\lesssim_{d,\kappa,\kappa_{0},M_{0},T}$ $\,\|f\|_{L^{p}((0,T)\times\Omega;H^{-1,q}(\mathbb{T}^{d};{\mathbb{R}}^{d}))}$
$+$ $\,\|g\|_{L^{p}((0,T)\times\Omega;L^{2}(\mathbb{T}^{d};\ell^{2}({\mathbb{N}};{\mathbb{R}}^{d})))}.$
The comments below Theorem A.1 extend to the above result. In particular, the LHS(A.15) can be replaced by
@@
$\|v\|_{L^{p}(\Omega;H^{\theta,p}(0,T;\mathbb{H}^{1-2\theta,q}(\mathbb{T}^{d})))}\text{ for $\theta\in[0,\tfrac{1}{2})$ provided $p>2$}.$
Moreover, due to [4, Proposition 3.10], the above estimates also hold if $v(\cdot,0)\in L^{p}_{\mathscr{F}_{0}}(\Omega;\mathbb{B}^{1-2/p}_{q,p}(\mathbb{T}^{d}))$ provided on the RHS(A.15) one also add $\|v(\cdot,0)\|_{L^{p}(\Omega;\mathbb{B}^{1-2/p}_{q,p}(\mathbb{T}^{d}))}$.
@@
The proof is an extension of the one of Theorem A.1. Since the deterministic component of the turbulent Stokes system (A.12) is more complicated than the one of (A.1), we perform the perturbation argument used in Theorem A.1 twice. In particular, in the first step, we ensure that the deterministic part has maximal $L^{p}$-regularity which will be used to reduce to the case $f\equiv 0$ of (A.12).
Step 1: (Maximal $L^{p}$-regularity – deterministic problem) There exist $p_{1}>2$ and constants $(K_{p})_{p\in[2,p_{1}]}$ such that, for all $p\in[2,p_{1}]$, $q\in[2,p]$ and $f\in L^{p}({\mathbb{R}}_{+};\mathbb{H}^{-1,q})$ the following Cauchy problem
@@
on $\mathbb{T}^{d}$, has a unique strong solution $w\in L^{p}({\mathbb{R}}_{+};\mathbb{H}^{1,q})$ satisfying
$\|w\|_{L^{p}({\mathbb{R}}_{+};H^{1,q})}\leq K_{p}\|f\|_{L^{p}({\mathbb{R}}_{+};H^{-1,q})}.$
@@
Substep 1a: (Analysis of constants – case $\xi_{n}\equiv 0$) For each $p>2$ there exists $C_{p}>0$ satisfying ${\lim_{p\downarrow 2}\sup_{2\leq r\leq p}C_{r}<1/\kappa_{0}}$ such that, for all $q\in[2,p]$ and $f\in L^{p}({\mathbb{R}}_{+};\mathbb{H}^{-1,q})$,
(A.17) $\|w\|_{L^{p}({\mathbb{R}}_{+};H^{1,q})}\leq C_{p}\|f\|_{L^{p}({\mathbb{R}}_{+};H^{-1,q})}$
where $w\in L^{p}({\mathbb{R}}_{+};\mathbb{H}^{1,q})$ is the unique strong solution to (A.16) with $\xi_{n}\equiv 0$. As in Step 1 of Theorem A.1, the existence of such $w$ follows from the boundedness $H^{\infty}$-calculus of the operator $w\mapsto-\Delta w+w$ on $(\mathbb{H}^{1,q^{\prime}})^{*}$ with domain $\mathbb{H}^{1,q}$. By interpolation and $\kappa_{0}<\kappa$, it remains to show the validity of (A.17) with $p=q=2$ and $C_{2}=1/\kappa$. To prove the latter, note that, by computing $\frac{\mathrm{d}}{\mathrm{d}t}\|w\|_{L^{2}}^{2}$ we obtain
@@
Substep 1b: Proof of the claim of Step 1. We prove the claim of Step 1 by (the deterministic version of) the method of continuity [5, Proposition 3.13]. Thus, for all $\lambda\in[0,1]$ consider, on $\mathbb{T}^{d}$,
$\partial_{t}w_{\lambda}=\kappa(\Delta w_{\lambda}-w_{\lambda})+\lambda Q_{\xi}w_{\lambda}+f,\qquad w_{\lambda}(0,\cdot)=0.$
As in the proof of Theorem A.1, it remains to prove an estimate for $w_{\lambda}$ in $L^{p}({\mathbb{R}}_{+};H^{1.q})$ with constant independent of $\lambda\in[0,1]$. Firstly, by (A.13), it follows that
@@
(A.19) $\lim_{q_{1}\downarrow 2}\sup_{2\leq q<q_{1}}D_{q_{1}}=\kappa_{0}.$
By interpolation, it is enough to prove that (A.18) holds with $C_{2}=\sqrt{2\kappa_{0}}$ whenever $q=2$. Note that, for all $w\in\mathbb{H}^{1}$,
$\|Q_{\xi}w\|_{(\mathbb{H}^{1})^{*}}$ $=\frac{1}{2}\sup_{\phi\in\mathbb{H}^{1}\,:\,\|\phi\|_{H^{1}}\leq 1}\,\sum_{n\geq 1}\int_{\mathbb{T}^{d}}\mathbb{Q}[(\xi_{n}\cdot\nabla)v]\cdot\mathbb{Q}[(\xi_{n}\cdot\nabla)w]\,\mathrm{d}x$
$\leq\frac{1}{2}\sup_{\phi\in\mathbb{H}^{1}\,:\,\|\phi\|_{H^{1}}\leq 1}\,\|(\mathbb{Q}[(\xi_{n}\cdot\nabla)w])_{n\geq 1}\|_{L^{2}(\ell^{2})}\|(\mathbb{Q}[(\xi_{n}\cdot\nabla)\phi])_{n\geq 1}\|_{L^{2}(\ell^{2})}$
$\leq\frac{1}{2}\sup_{\phi\in\mathbb{H}^{1}\,:\,\|\phi\|_{H^{1}}\leq 1}\,\|((\xi_{n}\cdot\nabla)w)_{n\geq 1}\|_{L^{2}(\ell^{2})}\|((\xi_{n}\cdot\nabla)\phi)_{n\geq 1}\|_{L^{2}(\ell^{2})}$
(A.14)${{\leq}}\kappa_{0}\|\nabla w\|_{L^{2}},$
where $(i)$ we used that $\|\mathbb{Q}\|_{\mathscr{L}(L^{2})}=1$ as $\mathbb{Q}$ is an ortogonal projection on $L^{2}$.
Now, similarly to Step 2 of Theorem A.1, one can prove an a-priori estimate for $\|w_{\lambda}\|_{L^{p}({\mathbb{R}}_{+};H^{1,q})}$ with constant independent of $\lambda$ provided $2\leq q\leq p$ is sufficiently small due to (A.19), $\kappa_{0}<\kappa$ and substep 1a.
Step 2: Conclusion. The proof of Theorem A.2 follows by applying again the method of continuity. Hence, for $\lambda\in[0,1]$, we consider
@@
on $\mathbb{T}^{d}$, where $f\in L^{p}({\mathbb{R}}_{+}\times\Omega;(\mathbb{H}^{1,q^{\prime}})^{*})$ and $g\in L^{p}({\mathbb{R}}_{+}\times\Omega;L^{q}(\ell^{2}))$ are given progressively measurable processes. By the method of continuity [5, Proposition 3.13], it is enough to prove the existence of $p_{0}>2$ and $C_{0}>0$ such that, for all $p\in[2,p_{0}]$, $q\in[2,p]$, $\lambda\in[0,1]$ and progressively measurable processes $f$ and $g=(g_{n})_{n\geq 1}$ as above,
(A.21) $\|v_{\lambda}\|_{L^{p}({\mathbb{R}}_{+}\times\Omega;H^{1,q})}\leq C_{p}\big{(}\|f\|_{L^{p}({\mathbb{R}}_{+}\times\Omega;(\mathbb{H}^{1,q^{\prime}})^{*})}+\|g\|_{L^{p}({\mathbb{R}}_{+}\times\Omega;L^{q}(\ell^{2}))}\big{)}$
where $v_{\lambda}\in L^{p}({\mathbb{R}}_{+}\times\Omega;H^{1,q})$ is a strong solution to (A.20). Without loss of generality, we may assume $p_{0}\leq p_{1}$ where $p_{1}$ is as in Step 1 of the current proof. As in Step 2 of Theorem A.2, due to Step 1, it is enough to prove (A.21) with $f\equiv 0$. We now again repeat the argument of Theorem A.1 by analyzing first the constant in the energy inequality and afterwards, we argue by perturbation.
Substep 2a: (Analysis of constants – case $\lambda=0$ and $f\equiv 0$) For each $p\in[2,p_{1}]$ there exists $K_{p}>0$ satisfying ${\lim_{p\downarrow 2}\sup_{2\leq r\leq p}K_{p}=1}$ and
(A.22) $(\kappa-\kappa_{0})\big{\|}(|v|^{2}+|\nabla v|^{2})^{1/2}\big{\|}_{L^{p}({\mathbb{R}}_{+}\times\Omega;L^{q})}^{2}$
$+\|(\mathbb{P}[(\xi_{n}\cdot\nabla)v])_{n\geq 1}\|_{L^{p}({\mathbb{R}}_{+}\times\Omega;H^{1,q})}^{2}$ $\leq K_{p}\|g\|_{L^{p}({\mathbb{R}}_{+}\times\Omega;L^{q}(\ell^{2}))}^{2}$
where $v\in L^{p}({\mathbb{R}}_{+}\times\Omega;H^{1,q})$ is the strong solution to (A.7) with $\lambda=0$ and $f\equiv 0$. The existence of such $v$ follows from Step 1 and the argument in [71, Theorem 3.9] for the $g$-term. The intuition behind the RHS(A.22) is that such a quantity in the case $p=q=2$ appears naturally in the energy balance with $K_{2}=1$. To see this, first note that, by combining the Itô formula [59, Theorem 4.2.5] and standard integration by parts argument, one has
@@
$\|(\mathbb{P}[(\xi_{n}\cdot\nabla)v])_{n\geq 1}\|_{H^{1}}^{2}$ $=\|((\xi_{n}\cdot\nabla)v)_{n\geq 1}\|_{H^{1}}^{2}-\|(\mathbb{Q}[(\xi_{n}\cdot\nabla)v])_{n\geq 1}\|_{H^{1}}^{2}$
(A.14)${{\leq}}\kappa_{0}\|\nabla v\|_{L^{2}}^{2}-\|(\mathbb{Q}[(\xi_{n}\cdot\nabla)v])_{n\geq 1}\|_{H^{1}}^{2}.$
Combing the above and (A.23), we get that (A.22) holds with $q=p=2$ and $K_{2}=1$.
We conclude by arguing by interpolation. To this end, consider the operator

### v3

@@ A.2. Stochastic Meyers’ estimates for the turbulent Stokes system
In this subsection, we prove the Meyers’ estimates for the turbulent Stokes system on $\mathbb{T}^{d}$, i.e.,
(A.16) $\left\{\begin{aligned} \partial_{t}v&=\kappa\Delta v+Q_{\xi}v+f+\sum_{n\geq 1}\big(\mathbb{P}\big[(\xi_{n}\cdot\nabla)v\big]+g_{n}\big)\,\dot{W}_{t}^{n}&\text{ on }&\mathbb{T}^{d},\\ v(0,\cdot)&=0&\text{ on }&\mathbb{T}^{d},\end{aligned}\right$.
@@
Here and below, we use the notation introduced at the beginning of Section 5. Although it holds that $(\mathbb{H}^{1})^{*}=\mathbb{H}^{-1}$, here we will not employ such identification as it is important to keep track of constants. Note that $Q_{\xi}$ corresponds to the Itô-correction for the Stratonovich formulation of the transport noise for the turbulent Stokes system, see Section 5 and [10, Section 1].
Strong solutions to (A.16) can be defined similarly to those of (A.1). The following complements [10, Theorem 3.2] in case of $L^{\infty}$-transport noise.
@@
Let $\kappa>0$. Assume that
$\xi_{n}=(\xi_{j}^{n})_{j=1}^{d}:{\mathbb{R}}_{+}\times\Omega\times\mathbb{T}^{d}\to{\mathbb{R}}^{d}\text{ is $\mathscr{P}\otimes\mathscr{B}(\mathbb{T}^{d})$-measurable for all $n\geq 1$},$
and that there exist $M_{0}>0$ and $\kappa_{0}\in[0,\kappa)$ such that, a.e. on ${\mathbb{R}}_{+}\times\Omega\times\mathbb{T}^{d}$,
(A.17) $\|(\xi_{n})_{n\geq 1}\|_{\ell^{2}}$ $\leq M_{0},$ $(\text{boundedness})$
(A.18) $\frac{1}{2}\sum_{n\geq 1}\big(\xi_{n}\cdot\eta\big)^{2}$ $\leq\kappa_{0}|\eta|^{2}\ \text{ for all }\eta\in{\mathbb{R}}^{d}.$ $(\text{parabolicity})$
Then there exists $p_{0}(d,\kappa,\kappa_{0},M_{0})>2$ such that, for all stopping times $\tau$ with values $[0,T]$ with $T<\infty$, $p\in[2,p_{0}]$, $q\in[2,p]$ and progressively measurable processes $f$ and $g$ satisfying
$f\in L^{p}((0,\tau)\times\Omega;\mathbb{H}^{-1,q}(\mathbb{T}^{d}))\quad\text{ and }\quad g=(g_{n})_{n\geq 1}\in L^{p}((0,\tau)\times\Omega;\mathbb{L}^{q}(\mathbb{T}^{d};\ell^{2})),$
there exists a unique strong solution $v\in L^{p}((0,\tau)\times\Omega;\mathbb{H}^{1,q}(\mathbb{T}^{d}))$ to (A.16) and
(A.19) $\|v\|_{L^{p}((0,\tau)\times\Omega;H^{1,q}(\mathbb{T}^{d};{\mathbb{R}}^{d}))}\lesssim_{d,\kappa,\kappa_{0},M_{0},T,p}$ $\,\|f\|_{L^{p}((0,\tau)\times\Omega;H^{-1,q}(\mathbb{T}^{d};{\mathbb{R}}^{d}))}$
$+$ $\,\|g\|_{L^{p}((0,\tau)\times\Omega;L^{q}(\mathbb{T}^{d};\ell^{2}({\mathbb{N}};{\mathbb{R}}^{d})))}.$
The comments below Theorem A.1 extend to the above result. In particular, if $\tau=\infty$, then the LHS(A.19) can be replaced by
@@
$\|v\|_{L^{p}(\Omega;H^{\theta,p}(0,T;\mathbb{H}^{1-2\theta,q}(\mathbb{T}^{d})))}\text{ for $\theta\in[0,\tfrac{1}{2})$ provided $p>2$},$
and the first of the above also holds with $p=2$ (and thus $q=2$). Moreover, due to [5, Proposition 3.10], the above estimates also hold if $v(\cdot,0)\in L^{p}_{\mathscr{F}_{0}}(\Omega;\mathbb{B}^{1-2/p}_{q,p}(\mathbb{T}^{d}))$ provided on the RHS(A.19) one also add $\|v(\cdot,0)\|_{L^{p}(\Omega;\mathbb{B}^{1-2/p}_{q,p}(\mathbb{T}^{d}))}$.
@@
The proof is a variant of the one used in Theorem A.1. In particular, we consider the following SPDE:
(A.20) $\left\{\begin{aligned} \partial_{t}v&=\kappa(\Delta v-v)+Q_{\xi}v+f+\sum_{n\geq 1}\big(\mathbb{P}\big[(\xi_{n}\cdot\nabla)v\big]+g_{n}\big)\,\dot{W}_{t}^{n}&\text{ on }&\mathbb{T}^{d},\\ v(0,\cdot)&=0&\text{ on }&\mathbb{T}^{d},\end{aligned}\right$.
which is introduced for technical convenience and allows us to obtain a constant $p_{0}>2$ that is independent of the time horizon $T<\infty$, as stated in Theorem A.3; see also Remark A.2. Thus, in the remaining part of the proof, we prove existence and the estimate (A.19) for strong solutions to (A.20).
Since the deterministic part of the turbulent Stokes system (A.20) (i.e., with $\xi_{n}\equiv 0$) is more intricate than that of (A.1), we apply the perturbation argument–via the method of continuity from Theorem A.1–in two stages. In the first step, we establish maximal $L^{p}$-regularity for the deterministic component, which then enables a reduction to the case $f\equiv 0$ in (A.20).
Step 1: (Maximal $L^{p}$-regularity – deterministic problem) There exist $p_{1}>2$ and constants $(K_{p})_{p\in[2,p_{1}]}$ such that, for all $p\in[2,p_{1}]$, $q\in[2,p]$, $T\in(0,\infty]$, and $f\in L^{p}(0,T;\mathbb{H}^{-1,q})$ the following Cauchy problem
@@
on $\mathbb{T}^{d}$, has a unique strong solution $w\in L^{p}(0,T;\mathbb{H}^{1,q})$ satisfying
$\|w\|_{L^{p}(0,T;H^{1,q})}\leq K_{p}\|f\|_{L^{p}(0,T;H^{-1,q})}.$
@@
Substep 1a: (Analysis of constants – case $\xi_{n}\equiv 0$) For each $p>2$ there exists $C_{p}>0$ satisfying ${\limsup_{p\downarrow 2}C_{p}<1/\kappa_{0}}$ such that, for all $q\in[2,p]$, $T\in(0,\infty]$ and $f\in L^{p}(0,T;\mathbb{H}^{-1,q})$, there exists a unique strong solution $w\in L^{p}(0,T;\mathbb{H}^{1,q})$ to (A.21) $\xi_{n}\equiv 0$ (and hence $Q_{\xi}\equiv 0$) and
(A.22) $\|w\|_{L^{p}(0,T;H^{1,q})}\leq C_{p}\|f\|_{L^{p}(0,T;H^{-1,q})}.$
As in Step 1 of Theorem A.1, the existence of such $w$ follows from the boundedness $H^{\infty}$-calculus of the operator $w\mapsto-\Delta w+w$ on $(\mathbb{H}^{1,q^{\prime}})^{*}$ with domain $\mathbb{H}^{1,q}$. By repeating the argument in (A.10)-(A.12) and $\kappa_{0}<\kappa$, it remains to show the validity of (A.22) with $p=q=2$ and $C_{2}=1/\kappa$. To prove the latter, note that, by computing $\frac{\mathrm{d}}{\mathrm{d}t}\|w\|_{L^{2}}^{2}$ we obtain
@@
Substep 1b: Proof of the claim of Step 1. We prove the claim of Step 1 by (the deterministic version of) the method of continuity [6, Proposition 3.13] (the reader is referred to the text at the beginning of Step 2 in Theorem A.1 for some comments). Thus, for all $\lambda\in[0,1]$ consider, on $\mathbb{T}^{d}$,
(A.23) $\partial_{t}w_{\lambda}=\kappa(\Delta w_{\lambda}-w_{\lambda})+\lambda Q_{\xi}w_{\lambda}+f,\qquad w_{\lambda}(0,\cdot)=0.$
As in the proof of Step 2 of Theorem A.1, from the method of continuity and Step 1 of the current proof, it remains to prove that any strong solution $w_{\lambda}$ to the above that lies in $L^{p}(0,T;H^{1.q})$ for some $T\in(0,\infty]$ satisfies an a priori estimate with constant independent of $\lambda\in[0,1]$. Firstly, by (A.17), it follows that
@@
(A.25) $\lim_{q\downarrow 2}D_{q}=\kappa_{0}.$
By interpolation (again, see (A.10)-(A.12)), it is enough to prove that (A.24) holds with $C_{2}=\sqrt{2\kappa_{0}}$ whenever $q=2$. Note that, for all $w\in\mathbb{H}^{1}$,
$\|Q_{\xi}w\|_{(\mathbb{H}^{1})^{*}}$ $=\frac{1}{2}\sup_{\phi\,\in\,\mathbb{H}^{1}\,:\,\|\phi\|_{H^{1}}\leq 1}\,\sum_{n\geq 1}\int_{\mathbb{T}^{d}}\mathbb{Q}[(\xi_{n}\cdot\nabla)w]\cdot\mathbb{Q}[(\xi_{n}\cdot\nabla)\phi]\,\mathrm{d}x$
$\stackrel{{\scriptstyle(i)}}{{\leq}}\frac{1}{2}\sup_{\phi\,\in\,\mathbb{H}^{1}\,:\,\|\phi\|_{H^{1}}\leq 1}\,\|(\mathbb{Q}[(\xi_{n}\cdot\nabla)w])_{n\geq 1}\|_{L^{2}(\ell^{2})}\|(\mathbb{Q}[(\xi_{n}\cdot\nabla)\phi])_{n\geq 1}\|_{L^{2}(\ell^{2})}$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}\frac{1}{2}\sup_{\phi\,\in\,\mathbb{H}^{1}\,:\,\|\phi\|_{H^{1}}\leq 1}\,\|((\xi_{n}\cdot\nabla)w)_{n\geq 1}\|_{L^{2}(\ell^{2})}\|((\xi_{n}\cdot\nabla)\phi)_{n\geq 1}\|_{L^{2}(\ell^{2})}$
@@
where in $(i)$ we used the Cauchy-Scwartz inequality in $L^{2}(\ell^{2})=L^{2}(\mathbb{T}^{d};\ell^{2}({\mathbb{N}};{\mathbb{R}}^{d}))$, and in $(ii)$ that $\|\mathbb{Q}\|_{\mathscr{L}(L^{2})}=1$ as $\mathbb{Q}$ is an ortogonal projection on $L^{2}$.
Now, similarly to Step 2 of Theorem A.1, one can prove an a-priori estimate for $\|w_{\lambda}\|_{L^{p}(0,T;H^{1,q})}$ with constant independent of $\lambda$ and $T\in(0,\infty]$ provided $2\leq q\leq p$ is sufficiently small due to (A.25), $\kappa_{0}<\kappa$ and substep 1a.
Step 2: Conclusion. The proof of Theorem A.3 follows by applying again the method of continuity. Hence, for $\lambda\in[0,1]$ and a stopping time $\tau$ taking values in $0,\infty]$, we consider
(A.26) $\left\{\begin{aligned} \partial_{t}v_{\lambda}&=\kappa(\Delta v_{\lambda}-v_{\lambda})+Q_{\xi}v_{\lambda}+f+\sum_{n\geq 1}\big(\lambda\,\mathbb{P}\big[(\xi_{n}\cdot\nabla)v_{\lambda}\big]+g_{n}\big)\,\dot{W}_{t}^{n},\\ v_{\lambda}(0,\cdot)&=0,\end{aligned}\right$.
on $\mathbb{T}^{d}$, where $f\in L^{p}((0,\tau)\times\Omega;(\mathbb{H}^{1,q^{\prime}})^{*})$ and $g\in L^{p}((0,\tau)\times\Omega;L^{q}(\ell^{2}))$ are given progressively measurable processes. By the method of continuity [6, Proposition 3.13], it is enough to prove the existence of $p_{0}>2$ and $C_{0}>0$ such that, for all $p\in[2,p_{0}]$, $q\in[2,p]$, $\lambda\in[0,1]$, and any stopping time $\tau$ and any progressively measurable processes $f$ and $g=(g_{n})_{n\geq 1}$ as above,
(A.27) $\|v_{\lambda}\|_{L^{p}((0,\tau)\times\Omega;H^{1,q})}\leq C_{p}\big(\|f\|_{L^{p}((0,\tau)\times\Omega;(\mathbb{H}^{1,q^{\prime}})^{*})}+\|g\|_{L^{p}((0,\tau)\times\Omega;L^{q}(\ell^{2}))}\big)$
where $v_{\lambda}\in L^{p}((0,\tau)\times\Omega;H^{1,q})$ is a given strong solution to (A.26). We emphasize that, as for (A.14), the key in the above estimate is the independence of $C_{p}$ on $\lambda\in[0,1]$. Moreover, due to the method of continuity, the solution $v_{\lambda}$ is already given a priori.
Next, we turn to the proof of (A.27) Without loss of generality, we may assume $p_{0}\leq p_{1}$ where $p_{1}$ is as in Step 1 of the current proof. As in Step 2 of Theorem A.1, due to Step 1, it is enough to prove (A.27) with $f\equiv 0$. We now repeat the argument of Theorem A.1 by analyzing first the constant in the energy inequality, and afterwards, we argue by perturbation.
Substep 2a: (Analysis of constants – case $\lambda=0$ and $f\equiv 0$) For each $p\in[2,p_{1}]$ there exists $K_{p}>0$ satisfying ${\limsup_{p\downarrow 2}K_{p}=1}$, such that for all stopping time $\tau$ with values in $0,\infty]$, any progressively measurable $g=(g_{n})_{n\geq 1}\in L^{p}((0,\tau)\times\Omega;L^{q}(\ell^{2}))$ there exists a unique strong solution $v\in L^{p}((0,\tau)\times\Omega;\mathbb{H}^{1,q})$ to (A.26) with $\lambda=0$ and $f\equiv 0$. Moreover, the latter satisfies
(A.28) $\Big\|\Big[(\kappa-\kappa_{0})(|v|^{2}+|\nabla v|^{2})+\|(\mathbb{P}[(\xi_{n}\cdot\nabla)v]\big)_{n\geq 1}\big\|^{2}_{\ell^{2}}\Big]^{1/2}\Big\|_{L^{p}((0,\tau)\times\Omega;L^{q})}$
$\leq K_{p}\|g\|_{L^{p}((0,\tau)\times\Omega;L^{q}(\ell^{2}))}$ .
In the following, $\tau$ is a fixed stopping time with values in $0,\infty]$. The existence of such $v$ follows from Step 1 and [78, Theorem 3.9] (or by approximation, the process $g$ with step processes and arguing as in (A.15)). In the remaining part of this step, we show (A.28) with the claimed limiting behaviour of $K_{p}$. The intuition behind the RHS(A.28) is that such a quantity in the case $p=q=2$ appears naturally in the energy balance with $K_{2}=1$. To see this, first note that, by combining the Itô formula [65, Theorem 4.2.5] and a standard integration by parts argument, one has
@@
$\|(\mathbb{P}[(\xi_{n}\cdot\nabla)v])_{n\geq 1}\|_{L^{2}(\ell^{2})}^{2}$ $=\|((\xi_{n}\cdot\nabla)v)_{n\geq 1}\|_{L^{2}(\ell^{2})}^{2}-\|(\mathbb{Q}[(\xi_{n}\cdot\nabla)v])_{n\geq 1}\|_{L^{2}(\ell^{2})}^{2}$
[…]
Combing the above with (A.29), one obtains (A.28) with $q=p=2$ and $K_{2}=1$.
We conclude by arguing by an interpolation argument similar to (A.10)-(A.12). To this end, consider the operator